Show by calculation what is the angle of deviation formed by plane mirror when light falls normally on the mirror

Grade:12

2 Answers

Rajat
5 years ago
Let the incident ray be AO and the reflected ray beo OB. Let the mirror be XY. Let the normal be ON. Angle of incidence (i) equals to ang. AON and angle of reflection (r) is ang. BON. extend the incident ray to OA'.
So the angle of deflection is equal to A'OB. A'OB= A'OY+BOY= AOX+ BOY ( vert opp angles) = (90-AON)+(90-BON)= 180-i-r 
in case the incident ray is normal to the mirror i=r=0. 
So angle of deflection = 180-0-0= 180.
All the best!!!!
Khimraj
5 years ago
When a ray of light falls normally on the surface of a plane mirror , the following things happen:
1. Incident ray is reflected back along the same path.
2. Angle if incidence is zero and angle of  reflection will be  zero [according to laws of reflection] .